# TurBLiMP: A Turkish Benchmark of Linguistic Minimal Pairs

### v1

@@ Abstract
Ezgi Başar  Francesca Padovani  Jaap Jumelet  Arianna Bisazza Center for Language and Cognition (CLCG), University of Groningen e.basar@student.rug.nl {f.padovani, j.w.d.jumelet, a.bisazza}@rug.nl
@@ 1 Introduction
A foundational insight in linguistics research is that applying minimal changes to a sentence can render it entirely acceptable or unacceptable to native speakers (Chomsky, 1965). Minimal pairs, as illustrated in Example (1), are a widely used diagnostic tool in linguistics.
@@
[…]
Minimal pairs have been a cornerstone of linguistic analysis for decades, and in recent years they have become a vital tool for the linguistic evaluation of language models (LMs). Warstadt et al. (2020) published the first large-scale English Benchmark of Linguistic Minimal Pairs (BLiMP) in an effort to systematically evaluate the linguistic knowledge of language models, and since then various benchmarks have been introduced for other languages.
We contribute to this growing collection by introducing the first Turkish benchmark of linguistic minimal pairs. TurBLiMP enriches the typological diversity of available linguistic evaluation benchmarks by incorporating a morphologically rich agglutinative language with highly flexible word order. While Turkish and other agglutinative languages like Finnish have been the object of several studies focusing on word-level morphology (Ismayilzada et al., 2025), the effects of word order flexibility and morphological complexity on the robustness of sentence-level grammatical judgments have not been studied in detail before. We fill this gap by introducing two sets of experimental minimal pair paradigms.
@@ 2 Minimal Pair Benchmarks
Minimal pairs have played an important role for evaluating the linguistic abilities of language models, targeting phenomena such as subject-verb agreement (Linzen et al., 2016), filler-gap dependencies (Wilcox et al., 2018), and negative polarity items (Jumelet and Hupkes, 2018). Warstadt et al. (2020) then established an English benchmark of 67,000 sentence pairs testing 67 paradigms through automated generation based on linguist-curated templates. This work inspired numerous adaptations for other languages, each employing different benchmark creation strategies. Benchmarks using a similar template-based approach as BLiMP include CLiMP (Chinese, Xiang et al., 2021), ZhoBLiMP (Chinese, Liu et al., 2024), BLiMP-NL (Dutch, Suijkerbuijk et al., 2025), and for Basque/Swahili/Hindi by Kryvosheieva and Levy (2025). Another approach is based on modifying Universal Dependency trees, which has been used by SLING (Chinese, Song et al., 2022), RuBLiMP (Russian, Taktasheva et al., 2024), and MultiBLiMP (Jumelet et al., 2025), a multilingual benchmark covering 101 languages. Other approaches include the extraction of minimal pairs from linguistics journals, employed by JBLiMP (Japanese, Someya and Oseki, 2023), manual creation of pairs, as done for Icelandic by Ármannsson et al. (2025), and the usage of LLMs for generating pairs, as done for Tamil and Indonesian by Leong et al. (2023).
Methodological innovations across these benchmarks reveal key trade-offs between scale, linguistic coverage, and data quality. Template-based generation enables large datasets but risks producing unnatural sentences (Vázquez Martínez et al., 2023), while manual extraction from literature or learner corpora ensures quality at the cost of scale. Some of the benchmarks incorporate hybrid approaches and human validation steps to balance these concerns. TurBLiMP too is the result of such hybrid approaches. While creating our benchmark, we developed strategies specifically adapted to the challenges of creating minimal pairs for Turkish.
@@ 3 Turkish Morphosyntax & NLP
Furthermore, Turkish has highly productive agglutinative morphology, whereby words typically consist of several morphemes attached to a root. Speakers can easily produce and understand numerous legitimate but low-frequency word forms through regular morphological processes, yielding substantially larger vocabulary requirements for LMs compared to analytic and fusional languages. Many syntactic phenomena are realized in Turkish through morphology, rather than by separate function words like in English and other Indo-European languages that form a large chunk of the world’s highest-resource languages. A salient example is subordination, which largely involves the use of suffixes to nominalize or adverbialize the verb of the embedded clause. For instance, the sentence I know that Elif likes Gaye translates to Elif’in Gaye’yi sevdiğini biliyorum whose structure can be intuitively conveyed as ‘I know the liking of Gaye by Elif’. Here, the nominalized verb ‘like’ takes an accusative case suffix as the object of ‘know’, but also a possessive agreement suffix corresponding to the genitive suffix taken by the subordinate subject ‘Elif’.
In general, agglutinative languages have been shown to be particularly challenging for neural models Gerz et al. (2018); Cotterell et al. (2016); Park et al. (2021); Arnett and Bergen (2025). Focusing on Turkish, Ataman et al. (2017) established that fixed vocabulary constraints combined with suboptimal sub-word segmentation significantly impair neural machine translation performance for agglutinative languages. Ismayilzada et al. (2025) studied LLMs’ ability to produce and systematically understand novel well-formed combinations of morphemes in Turkish and Finnish, and reported limited morphological generalization. These findings suggest that studying flexible-order, morphologically rich languages like Turkish can provide unique insights into the true linguistic capabilities of LMs beyond surface fluency.
@@ 4 TurBLiMP
Determiners Geçen hafta tad-ı damağ-ım-da kal-an [bir /$\varnothing$] tatlı ye-di-m. last week taste-acc palate-1sg.poss-loc stay-part [a /$\varnothing$] dessert eat-pst-1sg Last week, I ate a dessert with a taste that lingered on my tongue.
@@
NPI Licensing Kalabalığ-ın ön-ün-de [$\varnothing$ /*hiç] şarkı söyle-di-m. crowd-gen front-poss-loc [$\varnothing$ /*ever] song sing-pst-1sg I (*ever) sang in front of a crowd.
@@
Quantifiers Mağaza-da [$\varnothing$ /*çoğu] ayakkabı dene-di-m. store-loc [$\varnothing$ /*çoğu] shoe try_on-pst-1sg I tried on shoes in the store.
@@ 4.1 Phenomena
Principle B in Binding Theory (Chomsky, 1981) asserts that pronouns should be free in their binding domain, implying that pronouns should not refer to another entity in the same immediate clause. Unacceptable sentences are created by swapping an anaphora coreferring with the subject with a pronoun of similar features.
@@
While determiners are largely optional in Turkish, the indefinite article bir is sometimes required. When a direct object occurs immediately before the verb, its accusative case ending can be omitted. If such an object is modified by a relative clause, the indefinite article must precede the noun head (Arslan-Kechriotis, 2009). Unacceptable sentences in this phenomenon omit the obligatory determiner.
@@
This phenomenon deals with a specific type of ellipsis called backward gapping. For coordinated clauses in Turkish, it is possible to omit the verb in the first clause, leading to a gap which is resolved by the verb in the second clause. Turkish only licenses this if both clauses maintain parallel word order (Bozşahin, 2000). Acceptable sentences show the same subject-object order across clauses while unacceptable ones alternate their order.
@@
The aorist is an aspect/ mood marker with three allomorphs -r, -Ir (high vowel harmony), and -Ar (non-high vowel harmony). While monosyllabic verbs take -Ar, a specific subset of irregular verbs take -Ir (Nakipoğlu et al., 2023). Unacceptable sentences feature an incorrect -Ar form.
@@
We focus on a specific type of island constraint in which complex noun phrases are modified by a relative clause containing a wh-phrase. The occurrence of the wh-phrase is only permitted if the wh-phrase is not an adjunct (Çakır, 2016). Acceptable sentences contain argument wh-phrases like who or what, while unacceptable ones contain wh-adjuncts such as how or why.
@@
Turkish extensively uses a derivational process called nominalization, where verbal bases take suffixes (like -DIK, -mA, and others) to form noun phrases. A category of Turkish verbs only selects complement clauses with -DIK, while others only allow -mA (Kornfilt, 2003b). Correspondingly, minimal pairs contain verbs with the correct and incorrect nominalization suffixes.
@@
Turkish licenses the passivization of intransitive verbs via passive suffixes, creating impersonal (vs. personal) passives. While personal passives permit optional by-phrases to express agents, impersonal passives prohibit them (Özsoy, 2009). Thus, acceptable sentences omit by-phrases, while unacceptable ones include them.
@@
Turkish quantifiers such as her and çoğu can only occur with accusative-marked nouns (Enç, 1991). All minimal pairs for this phenomenon feature direct objects without accusative marking. Unacceptable sentences include a quantifier before the bare noun while acceptable sentences omit it.
@@
Turkish uses participle suffixes -DIK and -An to form object and subject relative clauses (Göksel and Kerslake, 2005). -DIK clauses feature genitive-possessive agreement. The subject takes genitive case and the verb carries possessive agreement. In subject relative clauses with -An, only the object (if present) is case-marked. Minimal pairs target an argument preceding the nominalized verb. Acceptability depends on whether this noun is inflected with a genitive or non-genitive case ending.
@@
Turkish shows word order flexibility and allows postverbal scrambling. This means that constituents can appear after the verb in certain contexts. However, local postverbal scrambling from an embedded clause is prohibited (Kornfilt, 2003a). Acceptable sentences position the object before the embedded verb while unacceptable sentences feature them in the opposite order.
@@
Suspended affixation refers to a phenomenon where a shared suffix applies to all conjuncts in a coordinated structure, rather than being repeated. Turkish does not allow suspended affixation for predicates inflected only with the past tense suffix -DI (Serova, 2019). Minimal pairs feature two coordinated past-tense clauses. Acceptable sentences inflect both verbs, while unacceptable ones omit inflection on the first.
@@ 4.2 Benchmark Creation
To enhance lexical diversity, we then adopted a semi-automated workflow in which a masked Turkish LM, BERTurk (Schweter, 2020) is used to suggest lexical replacements at random positions of each manually created sentence. We verified and adjusted each replacement manually to ensure acceptability. This process yielded 100 samples per phenomenon. In a final fully-automated augmentation step, BERTurk was used to generate a list of contextually appropriate words for replacement (e.g. woman or boy for girl). We use the Turkish morphology pipeline by Akın and Akın (2007) to inflect them with the same morphological features. At the end of this process, our 100 manually validated pairs increase to 1000 pairs per phenomenon. Our three-fold approach balanced scalability with linguistic precision, resulting in a robust benchmark for evaluating Turkish LMs.
@@ 4.3 Experimental Paradigms
We further assess the robustness of LMs’ syntactic abilities by focusing on two salient properties of Turkish: (i) word order flexibility and (ii) subordination through morphological processes, both discussed in Section 3. Word order variations provide a useful framework for testing the effect of word order biases on syntactic competence, extending the types of variations covered by the existing minimal pair benchmarks Linzen et al. (2016); Mueller et al. (2020). Subordination is a particularly interesting case to study the interplay between syntactic competence and morphological generalization, broadening the scope of current word-level evaluations Ismayilzada et al. (2025).
@@ 5 Human Acceptability Judgments
To validate our benchmark, we collected acceptability judgments from 30 native Turkish speakers using a 7-point Likert scale (1: completely unacceptable, 7: completely acceptable). While previous BLiMP variants rely on forced-choice tasks for data validation, BLiMP-NL (Suijkerbuijk et al., 2025) collects Likert scale responses to capture the gradient nature of acceptability judgments. We followed their approach to provide a benchmark that allows for fine-grained evaluation of model-human alignment. Our participant pool was mixed, comprising 17 linguistics students and 13 non-linguists. The study was carried out via an anonymous online survey. Appendix B includes a screenshot of survey instructions. Each participant rated 216 sentences spanning 16 linguistic phenomena as well as 20 experimental paradigms. 3 acceptable and 3 unacceptable sentences were included for each grammatical category, and the acceptability conditions were flipped between the two survey versions.
@@
Figure 1 reports average acceptability judgments for each phenomenon. Additional participant rating statistics are provided in Appendix C. The responses are first normalized by transforming Likert scores to z-scores. Overall, participants made clear distinctions between acceptable and unacceptable sentences. Some phenomena such as Island Effects, Passives, and Nominalization were less discriminable than others.
@@ 6 Experimental Setup
We employed the Goldfish series Chang et al. (2024), a series of causal LMs with fixed architecture trained on varying training data sizes (5MB, 10MB, 100MB, and 1000MB). Another monolingual model we used is BERTurk (Schweter, 2020), a 185M-parameter Turkish masked LM. With a vocabulary size of 128k, it is the only masked LM in our set of monolingual models. The largest monolingual model that we test is cosmosGPT (Kesgin et al., 2024), a 774M-parameter GPT-2-based model pretrained on Turkish web corpora and books.
@@
The evaluated multilingual models include Qwen 2.5 7B (Yang et al., 2025), Llama 3.1 8B (Meta, 2024), Aya Expanse 8B (Dang et al., 2024), Gemma 2 7B (Team et al., 2024), Gemma 3 4B and 12B (Team et al., 2025), as well as EuroLLM 9B (Martins et al., 2024). For a balanced comparison between the various models, we employed comparable parameter sizes ranging from 4B to 12B. Notably, Aya Expanse is the only instruction-tuned variant in our set of multilingual models, supporting 23 languages including Turkish. The Gemma series also boast multilinguality with Gemma 3 providing support for over 140 languages. EuroLLM prioritizes the coverage of European languages alongside a few others including Turkish.
As our evaluation metric for model performance, we computed entire-sequence log probabilities for acceptable and unacceptable sentences in each pair using the minicons library Misra (2022); Kauf and Ivanova (2023). Accuracy scores reflect the proportion of pairs where the model assigned a higher probability to the acceptable sentence. We also report Pearson’s correlation between human and model evaluations, calculated from the difference between average scores of acceptable and unacceptable sentences.

### v2

@@ Abstract
Ezgi Başar   Francesca Padovani   Jaap Jumelet   Arianna Bisazza Center for Language and Cognition (CLCG), University of Groningen {e.basar, f.padovani, j.w.d.jumelet, a.bisazza}@rug.nl
@@ 1 Introduction
A foundational insight in linguistics research is that applying minimal changes to a sentence can render it entirely acceptable or unacceptable to native speakers (Chomsky1965-CHOAOT-2). Minimal pairs, as illustrated in Example (1), are a widely used diagnostic tool in linguistics.
@@
$\ex[*]People in Istanbul loves cat$s.
Minimal pairs have been a cornerstone of linguistic analysis for decades, and in recent years they have become a vital tool for the linguistic evaluation of language models (LMs). warstadt-etal-2020-blimp-benchmark published the first large-scale English Benchmark of Linguistic Minimal Pairs (BLiMP) in an effort to systematically evaluate the linguistic knowledge of language models, and since then various benchmarks have been introduced for other languages.
We contribute to this growing collection by introducing the first Turkish benchmark of linguistic minimal pairs. TurBLiMP enriches the typological diversity of available linguistic evaluation benchmarks by incorporating a morphologically rich agglutinative language with highly flexible word order. While Turkish and other agglutinative languages like Finnish have been the object of several studies focusing on word-level morphology (ismayilzada-etal-2025-evaluating), the effects of word order flexibility and morphological complexity on the robustness of sentence-level grammatical judgments have not been studied in detail before. We fill this gap by introducing two sets of experimental minimal pair paradigms.
@@ 2 Minimal Pair Benchmarks
Minimal pairs have played an important role for evaluating the linguistic abilities of language models, targeting phenomena such as subject-verb agreement (linzen-etal-2016-assessing), filler-gap dependencies (wilcox-etal-2018-rnn), and negative polarity items (jumelet-hupkes-2018-language). warstadt-etal-2020-blimp-benchmark then established an English benchmark of 67,000 sentence pairs testing 67 paradigms through automated generation based on linguist-curated templates. This work inspired numerous adaptations for other languages, each employing different benchmark creation strategies. Benchmarks using a similar template-based approach as BLiMP include CLiMP (Chinese, xiang-etal-2021-climp), ZhoBLiMP (Chinese, liu2024zhoblimpsystematicassessmentlanguage), BLiMP-NL (Dutch, 10.1162/coli_a_00559), and for Basque/Swahili/Hindi by kryvosheieva-levy-2025-controlled. Another approach is based on modifying Universal Dependency trees, which has been used by SLING (Chinese, song-etal-2022-sling), RuBLiMP (Russian, taktasheva-etal-2024-rublimp), and MultiBLiMP (jumelet2025multiblimp10massivelymultilingual), a multilingual benchmark covering 101 languages. Other approaches include the extraction of minimal pairs from linguistics journals, employed by JBLiMP (Japanese, someya-oseki-2023-jblimp), manual creation of pairs, as done for Icelandic by armannsson-etal-2025-icelandic, and the usage of LLMs for generating pairs, as done for Tamil and Indonesian by leong2023bhasaholisticsoutheastasian.
Methodological innovations across these benchmarks reveal key trade-offs between scale, linguistic coverage, and data quality. Template-based generation enables large datasets but risks producing unnatural sentences (javier-vazquez-martinez-etal-2023-evaluating), while manual extraction from literature or learner corpora ensures quality at the cost of scale. Some of the benchmarks incorporate hybrid approaches and human validation steps to balance these concerns. TurBLiMP too is the result of such hybrid approaches. While creating our benchmark, we developed strategies specifically adapted to the challenges of creating minimal pairs for Turkish.
@@ 3 Turkish Morphosyntax & NLP
Furthermore, Turkish has highly productive agglutinative morphology, whereby words typically consist of several morphemes attached to a root. Speakers can easily produce and understand numerous legitimate but low-frequency word forms through regular morphological processes, yielding substantially larger vocabulary requirements for LMs compared to analytic and fusional languages. Many syntactic phenomena are realized in Turkish through morphology, rather than by separate function words like in English and other Indo-European languages that form a large chunk of the world’s highest-resource languages. A salient example is subordination, which largely involves the use of suffixes to nominalize or adverbialize the verb of the embedded clause.
[…]
Elif’in Gaye’yi sev-diğ-in-i bil-iyor-um.
Elif-3SG.GEN Gaye-ACC like-NMLZ-3SG.POSS-ACC know-PROG-1SG
[…]
For instance, the subordination structure in Example 3 can be intuitively conveyed as ‘I know the liking of Gaye by Elif’. Here, the nominalized verb ‘like’ takes an accusative case suffix as the object of ‘know’, but also a possessive agreement suffix corresponding to the genitive suffix taken by the subordinate subject ‘Elif’.
In general, agglutinative languages such as Finnish, Tamil, Basque, Indonesian or Japanese have been shown to be particularly challenging for neural models gerz-etal-2018-relation; cotterell-etal-2016-sigmorphon; park-etal-2021-morphology; arnett-bergen-2025-language. Focusing on Turkish, Ataman2017LinguisticallyMV established that fixed vocabulary constraints combined with suboptimal sub-word segmentation significantly impair neural machine translation performance for agglutinative languages. ismayilzada-etal-2025-evaluating studied LLMs’ ability to produce and systematically understand novel well-formed combinations of morphemes in Turkish and Finnish, and reported limited morphological generalization. These findings suggest that studying flexible-order, morphologically rich languages like Turkish can provide unique insights into the true linguistic capabilities of LMs beyond surface fluency.
@@ 4 TurBLiMP
Determiners Geçen hafta tad-ı damağ-ım-da kal-an [bir $\varnothing$] tatlı ye-di-m. last week taste-acc palate-1sg.poss-loc stay-part [a $\varnothing$] dessert eat-pst-1sg Last week, I ate a dessert with a taste that lingered on my tongue.
@@
NPI Licensing Kalabalığ-ın ön-ün-de $\varnothing$ /*hiç] şarkı söyle-di-m. crowd-gen front-poss-loc $\varnothing$ /*ever] song sing-pst-1sg I (*ever) sang in front of a crowd.
@@
Quantifiers Mağaza-da $\varnothing$ /*çoğu] ayakkabı dene-di-m. store-loc $\varnothing$ /*çoğu] shoe try_on-pst-1sg I tried on shoes in the store.
@@ 4.1 Phenomena
Principle B in Binding Theory (chomsky1981) asserts that pronouns should be free in their binding domain, implying that pronouns should not refer to another entity in the same immediate clause. Unacceptable sentences are created by swapping an anaphora coreferring with the subject with a pronoun of similar features.
@@
While determiners are largely optional in Turkish, the indefinite article bir is sometimes required. When a direct object occurs immediately before the verb, its accusative case ending can be omitted. If such an object is modified by a relative clause, the indefinite article must precede the noun head (arslan2009referentiality). Unacceptable sentences in this phenomenon omit the obligatory determiner.
@@
This phenomenon deals with a specific type of ellipsis called backward gapping. For coordinated clauses in Turkish, it is possible to omit the verb in the first clause, leading to a gap which is resolved by the verb in the second clause. Turkish only licenses this if both clauses maintain parallel word order (Bozsahin2000). Acceptable sentences show the same subject-object order across clauses while unacceptable ones alternate their order.
@@
The aorist is an aspect/ mood marker with three allomorphs -r, -Ir (high vowel harmony), and -Ar (non-high vowel harmony). While monosyllabic verbs take -Ar, a specific subset of irregular verbs take -Ir (NAKIPOĞLU_UZUNDAĞ_KETREZ_2023). Unacceptable sentences feature an incorrect -Ar form.
@@
We focus on a specific type of island constraint in which complex noun phrases are modified by a relative clause containing a wh-phrase. The occurrence of the wh-phrase is only permitted if the wh-phrase is not an adjunct (sinancakir2016). Acceptable sentences contain argument wh-phrases like who or what, while unacceptable ones contain wh-adjuncts such as how or why.
@@
Turkish extensively uses a derivational process called nominalization, where verbal bases take suffixes (like -DIK, -mA, and others) to form noun phrases. A category of Turkish verbs only selects complement clauses with -DIK, while others only allow -mA (Kornfilt+2003+129+216). Correspondingly, minimal pairs contain verbs with the correct and incorrect nominalization suffixes.
@@
Turkish licenses the passivization of intransitive verbs via passive suffixes, creating impersonal (vs. personal) passives. While personal passives permit optional by-phrases to express agents, impersonal passives prohibit them (ozsoy2009). Thus, acceptable sentences omit by-phrases, while unacceptable ones include them.
@@
Turkish quantifiers such as her and çoğu can only occur with accusative-marked nouns (a5cf88a7-e671-314d-be95-4caaea877f88). All minimal pairs for this phenomenon feature direct objects without accusative marking. Unacceptable sentences include a quantifier before the bare noun while acceptable sentences omit it.
@@
Turkish uses participle suffixes -DIK and -An to form object and subject relative clauses (göksel2005turkish). -DIK clauses feature genitive-possessive agreement. The subject takes genitive case and the verb carries possessive agreement. In subject relative clauses with -An, only the object (if present) is case-marked. Minimal pairs target an argument preceding the nominalized verb. Acceptability depends on whether this noun is inflected with a genitive or non-genitive case ending.
@@
Turkish shows word order flexibility and allows postverbal scrambling. This means that constituents can appear after the verb in certain contexts. However, local postverbal scrambling from an embedded clause is prohibited (kornfiltscrambling). Acceptable sentences position the object before the embedded verb while unacceptable sentences feature them in the opposite order.
@@
Suspended affixation refers to a phenomenon where a shared suffix applies to all conjuncts in a coordinated structure, rather than being repeated. Turkish does not allow suspended affixation for predicates inflected only with the past tense suffix -DI (kutay2019). Minimal pairs feature two coordinated past-tense clauses. Acceptable sentences inflect both verbs, while unacceptable ones omit inflection on the first.
@@ 4.2 Benchmark Creation
To enhance lexical diversity, we then adopted a semi-automated workflow in which a masked Turkish LM, BERTurk (stefan_schweter_2020_3770924) is used to suggest lexical replacements at random positions of each manually created sentence. We verified and adjusted each replacement manually to ensure acceptability. This process yielded 100 samples per phenomenon. In a final fully-automated augmentation step, BERTurk was used to generate a list of contextually appropriate words for replacement (e.g. woman or boy for girl). We use the Turkish morphology pipeline by akin2007zemberek to inflect them with the same morphological features when applicable. At the end of this process, our 100 manually validated pairs increase to 1000 pairs per phenomenon. Our three-fold approach balanced scalability with linguistic precision, resulting in a robust benchmark for evaluating Turkish LMs.
@@ 4.3 Experimental Paradigms
We further assess the robustness of LMs’ syntactic abilities by focusing on two salient properties of Turkish: (i) word order flexibility and (ii) subordination through morphological processes, both discussed in Section 3. Word order variations provide a useful framework for testing the effect of word order biases on syntactic competence, extending the types of variations covered by the existing minimal pair benchmarks linzen-etal-2016-assessing; mueller-etal-2020-cross. Subordination is a particularly interesting case to study the interplay between syntactic competence and morphological generalization, broadening the scope of current word-level evaluations ismayilzada-etal-2025-evaluating.
@@ 5 Human Acceptability Judgments
To validate our benchmark, we collected acceptability judgments from 30 native Turkish speakers using a 7-point Likert scale (1: completely unacceptable, 7: completely acceptable). While previous BLiMP variants rely on forced-choice tasks for data validation, BLiMP-NL (10.1162/coli_a_00559) collects Likert scale responses to capture the gradient nature of acceptability judgments. We followed their approach to provide a benchmark that allows for fine-grained evaluation of model-human alignment. Our participant pool was mixed, comprising 17 linguistics students and 13 non-linguists. The study was carried out via an anonymous online survey. Appendix B includes a screenshot of survey instructions. Each participant rated 216 sentences spanning 16 linguistic phenomena as well as 20 experimental paradigms. 3 acceptable and 3 unacceptable sentences were included for each grammatical category, and the acceptability conditions were flipped between the two survey versions.
@@
Figure 1 reports average acceptability judgments for each phenomenon. Additional participant rating statistics are provided in Appendix C. The responses are first normalized by transforming Likert scores to z-scores. We assume that every participant uses the scale slightly differently, and this step ensures comparability across participants.
Overall, our analysis reveals that participants made clear distinctions between acceptable and unacceptable sentences. The clear separation in z-scores between grammatical and ungrammatical constructions confirms that the targeted syntactic distinctions are perceptible to native speakers. We can also note that some phenomena such as Island Effects, Passives, and Nominalization were less discriminable than others.
@@ 6 Experimental Setup
We employed the Goldfish series chang-etal-2024-goldfish, a series of causal LMs with fixed architecture trained on varying training data sizes (5MB, 10MB, 100MB, and 1000MB). Another monolingual model we used is BERTurk (stefan_schweter_2020_3770924), a 185M-parameter Turkish masked LM. With a vocabulary size of 128k, it is the only masked LM in our set of monolingual models. The largest monolingual model that we test is cosmosGPT (kesgin2024introducing), a 774M-parameter GPT-2-based model pretrained on Turkish web corpora and books.
@@
The evaluated multilingual models include Qwen 2.5 7B (qwen2025qwen25technicalreport), Llama 3.1 8B (meta2024llama3.1), Aya Expanse 8B (dang2024ayaexpansecombiningresearch), Gemma 2 7B (gemmateam2024gemma2improvingopen), Gemma 3 4B and 12B (gemmateam2025gemma3technicalreport), as well as EuroLLM 9B (martins2024eurollmmultilinguallanguagemodels). For a balanced comparison between the various models, we employed comparable parameter sizes ranging from 4B to 12B. Notably, Aya Expanse is the only instruction-tuned variant in our set of multilingual models, supporting 23 languages including Turkish. The Gemma series also boast multilinguality with Gemma 3 providing support for over 140 languages. EuroLLM prioritizes the coverage of European languages alongside a few others including Turkish.
As our evaluation metric for model performance, we computed entire-sequence log probabilities for acceptable and unacceptable sentences in each pair using the minicons library misra2022minicons; kauf2023better. Accuracy scores reflect the proportion of pairs where the model assigned a higher probability to the acceptable sentence. We also report Pearson’s correlation between human and model evaluations, calculated from the difference between average scores of acceptable and unacceptable sentences.